# Masked Mixers for Language Generation and Retrieval

### v1

@@ Abstract
Attention mechanisms that confer selective focus on a strict subset of input elements are nearly ubiquitous in language models today. We posit there to be downside to the use of attention: most information present in the input is necessarily lost. In support of this idea we observe poor input representation accuracy in transformers, but find more accurate representation in what we term masked mixers which replace self-attention with masked convolutions. Applied to TinyStories the masked mixer learns causal language tasks more efficiently than early transformer implementations and somewhat less efficiently than optimized, current implementations. The most efficient learning algorithm observed for this dataset is a transformer-masked mixer hybrid, suggesting that these models learn in an orthogonal manner. We hypothesized that the information loss exhibited by transformers would be much more detrimental to retrieval than generation, and to test this we introduce an efficient training approach for retrieval models based on existing generative model embeddings. With this method, embeddings from masked mixers are found to result in far better summary-to-story retrieval compared to embeddings from transformers.
@@ 1 Introduction
Since the introduction of the transformer [1], many deep learning application domains have witnessed a rapid increase in the state-of-the-art as a result of incorporation of this architecture. Originally designed for language processing tasks such as machine translation, the transformer has been found to be well suited to increases in model parameters such that ever-larger models may be trained on ever-larger datasets with continued increases in various metrics of goodness [2; 3; 4; 5].
@@
In spite of this impressive success, it is less than clear how efficiently these models may be trained for various language tasks: many types of architectures experience increases in goodness with an increased parameter numbers and dataset sizes. One way to investigate this question indirectly is to observe how information passes from the input to various hidden layers of the model, which can be considered to be closely related to the question of how these hidden layers represent the model’s input. Current investigations on this topic often focus on the method by which transformers transfer information to and from different tokens (c.f. [6]), but we focus on a simpler question: how much rather than how information is passed from the input to various parts of the model. In particular, we focus on the information present in the model’s last hidden layer, as this is most commonly used for tasks such as causal modeling (before the language head is applied) and retrieval.
@@
The goal of any attention operation is to focus on some input elements at the expense of others. Transformers used for language generation consist of many sequential self-attention modules (separated by MLPs) each with multiple attention heads per module, however, such that the restricted focus inherent in each attention transformation can no longer necessarily be ascribed to the model as a whole. In particular, deeper model layers no longer attend to embeddings of tokens but representations of attention of embeddings of tokens. In light of observations that trained language models tend to require high numerical precision from a tiny subset of all weights and activations which are often large in absolute value [7; 8] we hypothesized that the inherent information-limiting characteristic of attention does translate to the transformer model as a whole. Conversely, we hypothesized that a model with attention replaced via linear transformations would not exhibit these informational limitations.
In this work we observe that indeed transformer model hidden layers do not contain enough information to allow for accurate input reconstruction via gradient-based input representation algorithms, whereas the introduced ‘masked mixer’ models with self-attention substituted for masked convolutions do. We then refine the architecture of the convolutions using representation accuracy as a guide. We then compare causal language training efficiencies and autoregressive outputs of this model to that obtained for transformers, finding that the two architecture types are generally similarly efficient for causal language model training despite their very different input representation accuracies. We conclude by presenting evidence for the idea that masked mixers are more suitable than transformers for language retrieval.
@@ 1.1 Related Work
The body of work on deep learning representation is large. Much work towards understanding transformers has revolved around explanations for how attention transformations are capable of adding, removing, and transferring information from one token or patch to another (see for example [6] in the context of one-headed attention). For vision models, methods for applying gradient descent in order to observe representations in a model’s hidden layers were pioneered by such studies as [9]. Notably the method of representation employed in this work is in some ways much more straightforward than there: instead of performing smoothing and gradient orientation procedures to gain a comprehendable representation, we perform vanilla gradient descent followed by pseudoinversion to recover tokens.
@@
The substitution of self-attention for convolutions in the transformer was introduced independently by [10] and [11] as the MLP-Mixer, a vision model designed primarily for image classification tasks. The authors of the former work noted that for a given fixed-sized image dataset the MLP-Mixer performed slightly worse than vision transformers, but do not directly compare their efficiencies for fixed compute training runs [10]. A language application of the MLP mixer based on Bloom-filtered inputs was introduced in [12] but was not designed for causal language modeling, and presents a number of architectural decisions that make this model unsuitable for autoregressive language modeling tasks that transformers are commonly trained for. To our knowledge there has not been any attempts to adapt the MLP-Mixer to causal language modeling, masked language modeling, or retrieval tasks.
@@
The finding that information often passes linearly between tokens in simplified transformers [6] and the observation that purely linear deep models (ie with no nonlinearities between matrix multiplications) learn in very non-linear fashions even if they lack full representational power of nonlinear models [13] suggest that simple inter-token linear transformations (for example, 1-dimensional convolutions) may capture much of the dynamics of more complex inter-token transformations, such as those in self-attention modules.
@@ 2 Accurate self and non-self token representation in Masked Mixers
How much information does a model contain about its input? We use a gradient descent-based approach to address this question for language models. With respect to information that can be used to uniquely identify a language input, the answer for transformers is not very much. We introduce the masked mixer and show that this model accurately represents its inputs before and after training, and even accurately represent a limited number of non-self tokens.
@@ 2.1 Input Representation Background
In this work we measure the information present in a model’s hidden layer representation of the input by attempting to recover the input using gradient descent on an initially random input, minimizing a chosen metric on the activations of that hidden layer as described in [15]. The goal is to invert a model such that information in the hidden layer’s activations representation is used to identify the input that resulted in those activations. The development of this approach as it relates to language models is presented in a series of posts starting with [16].
@@ 2.2 Masked mixer architecture
Elsewhere it was observed that large and otherwise quite capable language models (Llama-2 7b and 70b, GPT2, etc.) exhibit relatively inaccurate input representation for all but the shallowest few hidden layers [14]. It was also observed that vision MLP-mixers ([11; 10]) have superior input representation accuracy for non-self tokens compared to vision transformers [17], leading us to wonder if a mixer adapted to causal language modeling could retain this accurate input representation. We hypothesized that adapting an MLP-mixer to the process of language generation would give us a model suitable for autoregressive language generation and other tasks, but with much more accurate input representation abilities.
@@
where ‘f’ and ‘d’ are both equal to the number of tokens in the context window and ‘k’ is the 1-dimensional kernel size. A 1D convolution with a kernel size of 1 can also be implemented using ‘torch.linear’ on the sequence dimension in which the triangular weight masking is a more pithy ‘conv.weight.data = torch.tril(conv.weight)’ after reshaping, but experimentally this does not perform as well as the 1D convolutional kernel method above nor is it as flexible.
To see a diagrammatic interpretation of how these operations act to enforce causal language modeling, see Figure 2. The primary difference between these operations and the method used to CLM masking during during transformer training is that in that case a mask is usually applied to a self-attention module’s activations, whereas here the mask is applied to the weights directly.
@@
We call these models ‘Masked Mixers’ to emphasize 1) that the mask is an intrinsic part of the model for both training and inference and 2) these mixers in some cases no longer resemble MLPs, as the most performant versions use convolutions with non-unitary kernels or projected multi-headed kernels.
@@ 2.3 Masked mixers but not transformers exhibit accurate input representations
then the normalized Hamming metric is given in Equation (6). In words, this metric is the fraction of indices of the input where the generated representation’s token does not match the input’s token. The smaller the normalized Hamming distance, the more similar the input is to the model’s representation and the more information that representation carries. We measure before training as well as during and after 12 hours of RTX 3060 training on the first 2M examples of the TinyStories [20] dataset.
@@
As observed elsewhere for larger language models, small transformers designed for TinyStories modeling exhibit very poor input representation such that the information present in the last hidden layer is incapable of recovering very little of an input when (2) is applied (Figure 3), although this improves somewhat with training. Masked mixers exhibit near-perfect representation before training, and larger models retain this characteristic even after training (Figure 3).
Figure 3: Masked mixers exhibit more accurate input representation than transformers. All models are $n_{l}=8$ and all transformers are llama models with $n_{h}=32$. In d) the transformer is $d_{m}=256$ and mixer $d_{m}=512$.
@@
Masked mixers of a sufficient $d_{m}$ are in some sense biased towards accurate input representation, as untrained models nearly always exhibit perfect input representation regardless of their actual randomly initialized values. In this vein, transformers of a sufficient $d_{model}$ are biased against accurate input representation. It is interesting therefore to observe that masked mixer and transformer input representation abilities converge somewhat during training (Figure 3d), which suggests that for the TinyStories dataset there is some range of optimal input representation accuracies that is associated with minimization of all-next-token prediction during causal language modeling. This convergence is very slow, however, especially for larger models such that the model type’s initial bias determines most behavior after training as well (Figure 3c). If we scale up the compute used for training these two models such that approximately ten times the amount of samples are shown to the model (using much larger batches) we find that indeed the gap in the Hamming metric between masked mixer and transformer is substantially reduced (Figure S4).
3 Masked Mixer and Transformer learning efficiencies
@@
To test the training efficiency of the Masked Mixer compared to the Transformer, we employ the same relatively small dataset as for representation studies: the first 2M training examples of TinyStories with all evaluation examples with a Llama-2 tokenizer of size 4096 trained on that dataset. Models are trained using a correspondingly small fixed compute budget: 12 hours on an Nvidia RTX 3060, or 2.25 hours on a 4x Nvidia V100 cluster for the training process. The Llama -style transformer models have no special architectural or hyperparameter settings, and as this model has the same loss function on the language modeling head as the Masked Mixer (cross-entropy loss, loss masks on padding tokens) with the same hyperparameters (a constant context window size of 512, AdamW optimizer [21], learning rates, cross-entropy loss function etc.) we can directly compare the loss achieved after the fixed compute has been applied.
3.1 Representation accuracy for Masked Mixers roughly correlates with training efficiency
A rough correlation is observed between representational accuracy and training efficiency for masked mixers: the original expanded-style architecture becomes more efficient as the hidden dimension increases, but the ‘flat’ masked mixers with one convolutional layer between tokens which have better self- and non-self token representation are more efficient learners still (Figure 4, Table S2, S3). Both the increased efficiency with increased $d_{model}$ size as well as the increased efficiency of flat masked mixers relative to expanded ones correlate with increases in representation accuracy for these configurations (Figure 3). For Llama models the relationship between untrained or trained input accuracy and training efficiency is less clear (Table S1).
Figure 4: Flat mixers train more slightly more efficiently than expanded ones.
It should be noted that the expansion factors for expanded masked mixers are in some sense a superfluous hyperparameter when CLM maskind is applied. This is because any expansion factors not equal to one leads to rectangular inter-token convolutional weight matrices after reshaping, and triangular masking these matrices leads to unused rows or columns (see Figure S1 a graphical explanation of this phenomenon). This effectively makes expanded masked mixers with an expansion factor greater than one capable of viewing fewer samples per given time. But even when this effect is negated by considering losses at identical step number (rather than compute time) the expanded masked mixers are still slightly less efficient learners than ‘flat’ masked mixers. For expansion factors less than one, triangular causal language masking results in the loss of input token information (see Figure S2 for an example). Because of the less accurate non-self token representation observed in the last section as well as the slightly worse training efficiency, we do not investigate expanded mixers further in this work.
3.2 Masked Mixers are more efficient learners than non-optimized modern transformers
Flat masked mixers turn out to not only be more efficient learners than expanded mixers but they are also more efficient than modern transformers (Llama-2 architecture) with default architectural hyperparameters, varying the $d_{model}$ after previously optimizing other hyperparameters such as the number of layers per model and batch sizes (Figure (5, Table S1, Table S3). A comparison between flat masked mixer and transformer memory requirements for various context and $d_{m}$ sizes may be found in Table S10 and Table S11.
Figure 5: Left: Flat masked mixers are more efficient learners than modern transformers with a default attention head number, all with $b=16$ batch size except for $d_{m}=2048$ mixer and $d_{m}=1024$ llama which have $b=8$ to fit in memory. Right: Reducing the number of attention heads increases TinStories training efficiency substantially.
3.3 Optimized versions of modern transformers are somewhat more efficient learners than optimized masked mixers
Whilst optimizing the masked mixers and transformers in order to assess which model experiences a more efficient learning process, we focused on a number of hyperparameters: the hidden dimension $d_{model}$, the number of transformer or mixer block layers $n_{l}$, the AdamW learning rate, the number of attention heads and the batch size. For the Llama model we found that the most important single hyperparameter was the number of attention heads: decreasing this from 32 heads (the Llama 2 default) to 4 heads leads to a dramatic reduction in loss achieved per compute allocated, and somewhat surprisingly even on a per-step basis (Figure 5, Table S4). This fully optimized transformer learns more efficiently than $d_{m}$-optimized masked mixers using the 3060 compute node (Tables S3, S4) but less efficiently on the 4x V100 node (Tables S5, S6).
Amidst these investigations, one newly implemented optimization was found to result in a substantial boost in transformer training efficiency: the addition of Flash Attention 2 [22] led to optimized Llama models reaching training and test accuracies below those obtained for masked mixers (Table S7).
3.4 Masked mixer training efficiency with multiple heads or larger convolutional kernel sizes
Next we investigated whether increasing the number of inter-token parameters in the masked mixer would lead to more efficient training. As we had already found that flat mixers are more efficient TinyStories learners than expanded mixers, we started by testing the efficiency of masked mixers with parallel convolutions. We observed no benefits when using two parallel convolutions, however: the per-step loss is slightly lower but fewer optimizer steps are taken.
Increasing the convolutional kernel’s size leads to an increase in the number of inter-token parameters by a factor of the kernel size without increasing the depth of the inter-token transformations. A depiction of how a convolutional kernel of size 2 acts in the masked mixer is portrayed in Figure 6 for convenience. We found that masked mixers with a convolution size of 4 outperforms those with size 1 using the 3060 node (S3), but that there was little difference for the 4x V100 (Table S5).
Figure 6: Convolutional kernel size 2 masked mixer with ‘same’ style padding
We also investigated whether or not the use of multiple masked convolutional heads would result in greater training efficiency. To do this, we project each input dimension into a certain number of heads and apply a 1D masked convolution on each head, before concatenating the outputs and projecting back to match the dimension of $d_{model}$ before the feedforward layers are applied. The training efficiency of a two-headed mixer is very slightly superior to that of a single-headed mixer, and is on par with a mixer with convolutional kernel of size 4 (Table S6). Both for increases in kernel size as well as adding convolutional heads the per-step loss was lower than for the vanilla flat masked mixer, but as fewer steps were taken with fixed compute the overall losses remained similar.
Figure 7: Multiheaded masked mixer architecture
It may be wondered whether or not an increase in focus in the masked mixer could lead to increased learning efficiency. One may do this by Softmax-transforming the 1D convolutional weights before re-masking. Adding the Softmax turns out to leads to much worse training efficiency than if these weights are not transformed regardless of the number of steps taken, however, indicating that increasing the ‘focus’ of the masked convolution via this transformation does not benefit training (Table S5).
3.5 Transformer-mixer hybrids learn most efficiently
Observations of the very different nature by which transformers and mixers represent their inputs suggests that these models might learn in fundamentally different but perhaps complementary ways. If the learned structures from these models were indeed complementary in some respect, combining these architectures could allow the resulting model to achieve greater training efficiency than either transformer or mixer.
We find this is indeed the case: inserting a masked 1D convolution into the output of the transformer layer and adding a residual (here termed the ‘Transfixer’) results in drops in cross-entropy loss even for optimized versions of Llama as shown in Figure 8. We observe a drop in both test and validation loss for the 3060 platform as well as the 4x V100 (the latter having Flash Attention 2 implemented) with the transformer-mixer hybrid relative to the transformer as shown in Figure 8 and Table S7.
Figure 8: Transformer-mixer hybrid architecture and training efficiency
It is interesting to note that the placement of the convolutional layer in the transformer-mixer hybrid is important to this efficiency gain: if one instead places the masked convolution in parallel with the self-attention layer the resulting losses are identical to the transformer model without this convolution and if the masked convolution in placed before the self-attention layer the resulting training is somewhat less efficient than with no convolutional layer (Table S7).
3.6 Masked mixers are more efficient learners than early transformers
In some sense it is unsurprising that current transformer implementations are somewhat more efficient learners than masked mixers because transformers have seen a number of architectural improvements and compute optimizations since their introduction seven years ago. It is useful therefore to compare the learning efficiencies of early transformer implementations of all-next-token causal language modeling to masked mixers with the same training method, as these can be thought of as being similar in their developmental stage. We chose a model that was introduced before improvements such as Rotary Positional Encoding [23] and Flash Attention 2 to name a couple of notable improvements.
Specifically we tested the Hugging Face implementation of GPT [2], a CLM-style transformer model which was introduced shortly after the original transformer architecture. We find that GPT lags well behind the masked mixer’s training and validation loss values both for default as well as learning rate, batch size, and head number-optimized versions of the transformer (Figure 9, Table S6). These results are consistent across compute scales, suggesting a general gap in learning efficiency between these model architectures for this dataset.
Figure 9: GPT versus mixer TinyStory training efficiency

### v2

@@ Abstract
Attention mechanisms that confer selective focus on a strict subset of input elements are nearly ubiquitous in language models today. We posit there to be downside to the use of attention: most input information is lost. In support of this idea we observe poor input representation accuracy in transformers and more accurate representation in what we term masked mixers, which replace self-attention with masked convolutions. The masked mixer learns causal language modeling more efficiently than early transformer implementations and even outperforms optimized, current transformers when training on small (<512) but not larger context windows. Evidence is presented for the hypothesis that differences in transformer and masked mixer training efficiencies for various tasks are best predicted by input representation accuracy, or equivalently global invertibility. We hypothesize that the information loss exhibited by transformers would be more detrimental to retrieval than generation, as the former is more closely approximated by a bijective and thus invertible function. We find that masked mixers are more effective retrieval models both when the pretrained embedding model is unchanged as well as when the embedding model is modified via cosine similarity-based InfoNCE loss minimization. A small masked mixer is shown to outperform a large and near state-of-the-art transformer-based retrieval model, despite the latter being trained with many orders of magnitude more data and compute.
Model $n=32$ 64 128 256 512 1024 2048 4096 8192
Transformer 95.0 92.3 88.9 84.5 79.1 72.6 65.7 57.6 50.4
e5 Mistral Instruct 95.1 93.3 90.9 88.5 85.6 82.2 78.4 73.5 68.8
Masked Mixer 98.2 97.2 95.8 93.4 90.2 86.5 81.7 76.6 70.6
Table 1: FineMath retrieval top-1 accuracy (%) for $n$ sample sizes, each denoting retrieval with one query, one matching target and $n-2$ nonmatching targets.
@@ 1 Introduction
Since the introduction of the transformer [1], many deep learning application domains have witnessed a rapid increase in the state-of-the-art as a result of incorporation of this architecture. Originally designed for language processing tasks such as machine translation, the transformer has been found to be well suited to increases in model parameters such that ever-larger models may be trained on ever-larger datasets with continued increases in various metrics of goodness [2; 3; 4; 5]. The use of attention without recurrent mechanisms proved to be extremely influential in the sequence modeling field and beyond and has led to the introduction of other modules that attempt to address constraints inherent in dot-product attention [6].
@@
In spite of this impressive success, it is less clear how efficiently transformers or other models using attention may be trained for various language tasks: many types of architectures experience increases in goodness with an increased parameter numbers and dataset sizes, making early claims that the transformer scaled in superior ways to other models somewhat dubious. One way to investigate this question indirectly is to observe how information passes from the input to various hidden layers of the model, which can be considered to be closely related to the question of how these hidden layers represent the model’s input. Current investigations on this topic often focus on the method by which transformers transfer information to and from different tokens (c.f. [7]), but we focus on a simpler question: how much rather than how information is passed from the input to various parts of the model. In particular, we focus on the information present in the model’s last hidden layer, as this is most commonly used for tasks such as causal modeling (with a language modeling head applied) and retrieval.
The goal of any attention operation is to focus on some input elements at the expense of others. Here we use the term ‘focus’ to attempt to avoid overloading ‘attention’ with too many meanings, and define the term as analogous to ‘weight’ or the amount a change in an input element affects a change in the output. Equivalently, any effective attention transformation is by definition not globally invertible because out-of-focus input elements may be modified without a change in the output.
@@
Transformers used for language generation consist of many sequential self-attention modules (separated by MLPs) each with multiple attention heads per module, however, such that the restricted focus inherent in each attention transformation can no longer necessarily be ascribed to the model as a whole. In light of observations that trained language models tend to require high numerical precision from a tiny subset of all weights and activations which are often large in absolute value [8; 9] we hypothesize that the inherent information-limiting characteristic of attention does translate to the transformer model as a whole. Conversely, we hypothesize that a model substituting linear transformations for attention would not exhibit these informational limitations.
In this work we observe that transformer model hidden layers do not contain enough information to allow for accurate input reconstruction via gradient-based input representation algorithms, whereas the introduced ‘masked mixer’ models with self-attention substituted for masked convolutions do. We then compare causal language training efficiencies for a number of modeling tasks and find that the more a task requires something similar to what would be termed model invertibility, the more the masked mixer tends to outperform the transformer for both small and larger datasets. In general we corroborate the findings of the original vision-based mixers which that find that attention is superfluous to much of the current model performance observed [10; 11].
We investigate the performance of masked mixers and transformers for language retrieval task where a query to text matching task. We introduce a novel approach to training an accessory retrieval model using the embeddings of a CLM-pretrained model, and find that the masked mixer’s embeddings result in far more accurate retrieval. We then investigate the ability of masked mixers and transformers to perform retrieval when the embedding model is trained directly via InfoNCE-based cosine similarity loss, and find that once again masked mixers are substantially more accurate than transformers in this paradigm. We conclude with the observation that a masked mixer trained using these methods exhibits more accurate retrieval than a near state-of-the-art transformer model trained using the same techniques but with approximately 50,000x the pretraining compute to provide more context into the scale of the difference between masked mixer and transformer retrieval abilities.
@@ 1.1 Related Work
The body of work on deep learning representation is large. Much work towards understanding transformers has revolved around explanations for how attention transformations are capable of adding, removing, and transferring information from one token or patch to another (see for example [7] in the context of one-headed attention). For vision models, methods for applying gradient descent in order to observe representations in a model’s hidden layers were pioneered by such studies as [12]. Notably the method of representation employed in this work is in some ways much more straightforward than there: instead of performing smoothing and gradient orientation procedures to gain a comprehensible representation, we perform vanilla gradient descent followed by pseudoinversion to recover tokens.
@@
The substitution of self-attention for convolutions in the transformer was introduced independently by [10] and [11] as the MLP-Mixer, a vision model designed primarily for image classification tasks. The authors of the former work noted that for a given fixed-sized image dataset the MLP-Mixer performed slightly worse than vision transformers, but do not directly compare their efficiencies for fixed compute training runs [10]. We modify the mixer architecture introduced in these works in a few notable ways for causal language modeling: we apply triangular masking directly to the convolutional (ie the sequence MLP) weight parameter matrices both during training and inference, replace the mixer layer’s two sequential linear transformations separated by an activation layer on the sequence dimension with a single masked convolution with no activation, and optionally use non-unitary convolutional kernels or multi-headed convolutions.
A language application of the MLP mixer based on Bloom-filtered inputs was introduced in [13] but was not designed for causal language modeling, and presents a number of architectural decisions that make this model unsuitable for autoregressive language modeling tasks that transformers are commonly trained for.
Elaborations on the use of convolutions in place of attention have met with success in attempts to model language with sub-quadratic complexity in the sequence dimension. In [14], Fu and colleagues replace attention with Monarch matrix-based sequence transformations for efficient sub-quadratic language modeling, but motivated by that goal the authors use a Monarch matrix-specific causal parametrization rather than parameter masking for causal language modeling. Poli and colleagues [15] introduced the Hyena architecture consisting of a recurrent structure with long convolutions combined with a gating operation with the same goal, and also do not perform direct masking. We comment that these works provide valuable insight into achieving near-linear complexity language modeling via efficient convolutional operations, but our approach to efficient language modeling focuses on global invertibility predicting training efficiency rather than sequence complexity and thus we proceed with an entirely different theoretical framework and do not investigate sub-quadratic complexity convolutional operations (Monarch or Toeplitz weight matrices etc.) which are by nature relatively inefficient to train with respect to loss achieved per compute quantity applied.
@@
The finding that information often passes linearly between tokens in simplified transformers [7] and the observation that purely linear deep models (ie with no nonlinearities between matrix multiplications) exhibit non-linear learning characteristics even if they lack the representational power of nonlinear models [16] suggest that simple inter-token linear transformations (for example, 1-dimensional convolutions) may capture much of the dynamics of more complex inter-token transformations, such as those in self-attention modules.
To appreciate the motivation behind this work it is vital to understand that accurate input representation is equivalent to global invertibility only within the feasible input domain, not any arbitrary input. This is important to appreciate because it explains how models that are composed of non-invertible transformations (such as every linear layer in this work) can exhibit global invertibility, and is significant departure from work focused on invertible model transformations such a attention [17].
2 Our contribution
We introduce the following:
  1. 1.
    An input representation method for language models to test for global invertibility
  2. 2.
    The masked mixer, a modification of the MLP mixer for language modeling along with an inference method
  3. 3.
    A bidirectional modeling implementation with higher throughput than BERT-style masked language modeling
  4. 4.
    An non-trivial autoencoding method for sequence models with excellent retrieval characteristics
  5. 5.
    A retrieval training method in which an accessory model is trained to match embeddings
  6. 6.
    Masked mixer adaptations for cosine similarity-based InfoNCE retrieval training
The primary insight behind this work is that differences in a model’s input representation accuracy, or equivalently its global invertibility, are capable of predicting a wide range of differences in training efficiency and performance between transformers and models of identical architecture except for the substitution of attention for linear transformations. Using this principle as a guide, we introduce a new foundational model architecture that far surpasses the performance of a modern and optimized transformer for retrieval.
@@ 3 Accurate self and non-self token representation in Masked Mixers
How much input information does a model contain in its hidden layers? More specifically, given some hidden layer values $O_{l}(a,\theta)$ can we regenerate the model’s input $a$ or in other words can we effectively invert the model? We use a gradient descent-based approach to address this question for language models. With respect to information that can be used to uniquely identify a language input, the answer for transformers is not very much. We introduce the masked mixer and show that this model accurately represents its inputs before and after training, and even accurately represent a limited number of non-self tokens.
Properly speaking, autoregressive language models such as transformers exist as multifunctions during training as there are many input elements and many output elements, but are used as (many-to-one) functions during inference as there is one next token output generated per forward pass. We investigate invertibility in both settings.
@@ 3.1 Input Representation Background
In this work we measure the information present in a model’s hidden layer representation of the input by attempting to recover the input using gradient descent on an initially random input, minimizing a chosen metric on the activations of that hidden layer as described in [18]. The goal is to invert a model such that information in the hidden layer’s activations representation is used to identify the input that resulted in those activations.
@@ 3.2 Masked mixer architecture
Elsewhere it was observed that large and otherwise capable language models (Llama-2 7b and 70b etc.) exhibit relatively inaccurate input representation for all but the shallowest few hidden layers [19]. It was also observed that vision MLP mixer models ([11; 10]) have superior input representation accuracy for non-self tokens compared to vision transformers [20], leading us to wonder if a mixer adapted to causal language modeling could retain this accurate input representation. We hypothesized that adapting an MLP-mixer to the process of language generation would give us a model suitable for autoregressive language generation and other tasks, but with much more accurate input representation abilities.
@@
where ‘f’ and ‘d’ are both equal to the number of tokens in the context window and ‘k’ is the 1-dimensional kernel size. A 1D convolution with a kernel size of 1 can also be implemented using ‘torch.linear’ on the sequence dimension in which the triangular weight masking is a more pithy ‘conv.weight.data = torch.tril(conv.weight)’ after reshaping, but experimentally this does not exhibit the throughput as well as the 1D convolutional kernel method above nor is it as flexible.
A diagrammatic interpretation of how these operations act to enforce causal language modeling is shown in Figure 2. The primary difference between these operations and the method used to CLM masking during during transformer training is that in that case a mask is usually applied to a self-attention module’s activations, whereas here the mask is applied to the weights directly.
@@
We call these models ‘Masked Mixers’ to emphasize 1) that the mask is an intrinsic part of the model for both training and inference and 2) these mixers in some cases no longer resemble MLPs, as they contain convolutions with non-unitary kernels or projected multi-headed kernels.
@@ 3.3 Masked mixers but not transformers exhibit accurate input representations
We now turn to the task of measuring input representation accuracy, and we introduce a method for achieving this goal. Notably, we observe input representation before and after causal language model training, because the training procedure can drastically change a model’s representation accuracy.
@@
then the normalized Hamming metric is given in Equation (6). In words, this metric is the fraction of indices of the input where the generated representation’s token does not match the input’s token while ignoring the input’s pad token positions. The smaller the normalized Hamming distance, the more similar the input is to the model’s representation and the more information that representation carries. We measure before training as well as during and after 12 hours of RTX 3060 training on the first 2M examples of the TinyStories [23] dataset.
@@
As observed elsewhere for larger language models, small transformers designed for TinyStories modeling exhibit very poor input representation such that the information present in the last hidden layer is incapable of recovering very little of an input when (2) is applied (Figure 3), although this improves somewhat with training. Masked mixers exhibit near-perfect representation before training, and larger models retain this characteristic even after training on Tinystories (Figure 3).
Figure 3: Masked mixers exhibit more accurate input representation than transformers. All models are $n_{l}=8$ and all transformers are llama models with the default $n_{h}=32$. In d) the transformer is $d_{m}=256$ and mixer $d_{m}=512$.
@@
Masked mixers of a sufficient $d_{m}$ are in some sense biased towards accurate input representation, as untrained models nearly always exhibit perfect input representation regardless of their actual randomly initialized values. In this vein, transformers of a sufficient $d_{model}$ are biased against accurate input representation. It is interesting therefore to observe that masked mixer and transformer input representation abilities converge somewhat during training (Figure 3d), which suggests that for the TinyStories dataset there is some range of optimal input representation accuracies that is associated with minimization of all-next-token prediction during causal language modeling. This convergence is very slow, however, especially for larger models such that the model type’s initial bias determines most behavior after training as well (Figure 3c).
If we scale up the compute to increase the number of TinyStories samples by a factor of 10x we find that indeed the gap in the Hamming metric between masked mixer and transformer is substantially reduced (Figure S4), as is true when training is performed on a much larger and more diverse dataset with >10x the compute (Figure S5). Notably, masked mixers trained on those large datasets experience a general decrease in representational accuracy as the context window size of the training inputs increases (Figure S5).
4 Masked Mixer and Transformer Causal Language Modeling Efficiency
@@
To test the training efficiency of the Masked Mixer compared to the Transformer, we employ the same two datasets as for representation studies: the first train using 2M training examples of TinyStories with all evaluation examples with a Llama-2 tokenizer of size 4096 trained on that dataset, where models are trained using a correspondingly small fixed compute budget of 12 hours on an Nvidia RTX 3060, or 2.25 hours on a 4x Nvidia V100 cluster. We then train on a much larger and more varied dataset, specifically a 10 billion token subset of a extensively refined of recent Common Crawl snapshots called the FineWeb-edu [24] (hereafter termed ‘FineWeb’ for simplicity) with approximately ten times the compute amount, or approximately 20 hours on a 4x V100 server, with 200k steps taken per training run.
As the Llama -style transformer uses the same Cross-Entropy Loss function (with loss masking on pad token outputs) on the language modeling head as the Masked Mixer with the same hyperparameters (a context window of 512 unless otherwise noted, AdamW optimizer [25], etc.), we can directly compare the loss achieved after the fixed compute quantity has been applied.
4.1 Architecture Optimizations
We perform extensive architectural experiments to optimize the masked mixer’s architecture using TinyStories, and found that the use of input representation is an effective guide to training efficiencies for this model. Most notably, we modify the original MLP Mixer architecture to use only one convolution and no nonlinear activation between sequence elements as this was found to exhibit superior input representation accuracy and more efficient training (c.f. Section 9.1). The number of layers, hidden layer width, batch sizes, and hyperparameters like learning rates (all with linear decay) were then optimized on the same dataset via independent line searches. We found that multi-headed masked mixers and masked mixers with non-unitary convolutional kernel sizes were no more efficient, and Softmax-transformed convolutional weight mixers were substantially less efficient, than the standard one-masked-convolution architecture for TinyStories CLM training (Section 9.2).
We found that the optimized masked mixer exhibited more efficient TinyStories training characteristics than non-optimized Llama-style transformers (Section 9.3) but considered this to be an unfair comparison and proceeded to optimize the transformer model. Most notably here, we found that the default self-attention head number was much too large for optimal training on this dataset and reduced the default (32 heads) to four (Section 9.4). As for the masked mixer, we also optimized the number of hidden layers, layer widths, batch sizes, and learning rates using the same dataset. Once this was performed, it was found that the optimized Llama-style transformer was somewhat more efficient for CLM training on TinyStories than the optimized masked mixers.
As a general note, most comparisons between masked mixers and transformers in this work are performed between models with the same number of layers (usually $n_{l}=8$ or $n_{l}=16$) but with $d_{m}=512$ layer width for transformers and $d_{m}=1024$ for masked mixers. This difference in $d_{m}$ is the result of the optimizations detailed above: masked mixers require much less device memory as they contain fewer inter-token parameters (Tables S8 and S9), and as such these models are approximately equivalent with respect to on-device memory during training. During the aforementioned optimizations, we find empirically that increasing the transformer’s width to $d_{m}=1024$ with a decrease in batch size results in worse training efficiency and likewise decreasing the masked mixer’s width to $d_{m}=512$ while increasing the batch size leads to worse efficiency. We consider this model-dependent optimization as vital to valid CLM efficiency comparisons, as otherwise one can simply choose architectural or hyperparameter constraints that favor one or the other model. We open our optimization results open to the public in this work and urge other researchers to do the same, as the failure to perform at least some reasonable amount of architectural or hyperparameter optimization leads to relatively meaningless comparisons between architectures.